# FasterCache: Training-Free Video Diffusion Model Acceleration with High Quality

### v1

@@ A.3 Additional Qualitative Experiments
More visual results on Text-to-Video models The additional visual comparison results for Open-Sora 1.2 (Zheng et al., 2024), Open-Sora-Plan (Lab & etc., 2024), and Latte (Ma et al., 2024a) are presented in Fig. 13, Fig. 14, and Fig. 15, while further comparisons for CogVideoX (Yang et al., 2024) and Vchitect-2.0 Vchitect (2024) are shown in Fig. 16. Our method demonstrates reliable fidelity across various models and styles or content in video synthesis, while simultaneously achieving acceleration.
@@
Figure 16: More visual results on CogVideoX (480P 48 frames) & Vchitect-2.0 (480P 40frames).
@@
A.4 Additional Quantitative Experiments
User preference study To assess the effectiveness of our FasterCache, we additionally conduct a human evaluation. We randomly selected 30 videos for each model. Each rater receives a text prompt and two generated videos from different sampling acceleration methods (in random order). They are then asked to select the video with better visual quality. Five raters evaluate each sample, and the voting results are summarized in Table 5. As one can see, compared to other acceleration methods, the raters strongly prefer the videos generated by our method.

### v2

@@ A.4 Additional Qualitative Experiments
More visual results on Text-to-Video models The additional visual comparison results for Open-Sora 1.2 (Zheng et al., 2024), Open-Sora-Plan (PKU-Yuan Lab and Tuzhan AI etc., 2024), and Latte (Ma et al., 2024a) are presented in Fig. 17, Fig. 18, and Fig. 19, while further comparisons for CogVideoX-2B (Yang et al., 2024) and Vchitect-2.0 (Fan et al., 2025) are shown in Fig. 20. Our method demonstrates reliable fidelity across various models and styles or content in video synthesis, while simultaneously achieving acceleration.
Additionally, Fig. 21 demonstrates the visual performance of FasterCache on state-of-the-art models CogVideoX-5B and Mochi-10B (Team, 2024). FasterCache achieves an acceleration of 1.63 times (206s $\rightarrow$ 126s) on CogVideoX-5B and 1.74 times (320s $\rightarrow$ 184s) on Mochi-10B. As model scale increases, FasterCache consistently accelerates the sampling process while maintaining fidelity in synthesized videos. We also observe that as the generative capability of the base model improves, FasterCache becomes more robust in synthesizing videos with complex scenes or rapid motion. For instance, in Fig. 21, the $1st$ example shows subtle details of small groups of fish, the $3rd$ example highlights intricate finger details and complex non-rigid motions, and the $4th$ and $5th$ examples exhibit rapid and large-scale movements. These results demonstrate the broad potential of FasterCache in practical applications.
@@
Figure 20: More visual results on CogVideoX-2B (480P 48 frames) & Vchitect-2.0 (480P 40frames).
Figure 21: More visual results on CogVideoX-5B and Mochi-10B. Zoom in for details.
@@
A.5 Additional Quantitative Experiments
A.5.1 User preference study
To assess the effectiveness of our FasterCache, we additionally conduct a human evaluation. We randomly selected 30 videos for each model. Each rater receives a text prompt and two generated videos from different sampling acceleration methods (in random order). They are then asked to select the video with better visual quality. Five raters evaluate each sample, and the voting results are summarized in Table 6. As one can see, compared to other acceleration methods, the raters strongly prefer the videos generated by our method.
@@ A.5.1 User preference study
A.5.2 Hyperparameter Selection
Table 7: Different Dynamic FR caching intervals.
Interval LPIPS PSNR SSIM
2 0.0590 28.41 0.8938
3 0.0698 27.95 0.8853
4 0.0751 27.61 0.8823
5 0.0897 27.39 0.8712
Table 8: Different CFG-Cache caching intervals.
Interval LPIPS PSNR SSIM
1 0.0496 28.88 0.8964
3 0.0537 28.56 0.8947
5 0.0590 28.41 0.8938
7 0.0724 27.68 0.8818
9 0.0104 27.44 0.8706
Caching timestep interval of Dynamic Feature Reuse We experimented with different caching timestep intervals for Dynamic Feature Reuse. According to Table 8, it can be observed that as the caching timestep interval increases, the fidelity of the synthesized results gradually decreases. In practice, the caching timestep interval for Dynamic Feature Reuse can be adjusted as needed.
Caching timestep interval of CFG-Cache We experimented with different CFG-Cache intervals and found that when the interval exceeds 5 timesteps, there is a significant decline in fidelity, as shown in Table 8. Therefore, to balance fidelity and efficiency, we chose a CFG-Cache caching interval of 5. This means that after CFG-Cache is initiated, the model performs full inference for both the conditional and unconditional branches every 5 timesteps and caches the features.
The configuration of $\alpha$ in CFG-Cache. In CFG-Cache, we experimented with different configurations of $\alpha$, where $\alpha_{1}$ is used to enhance low-frequency biases and $\alpha_{2}$ is used to enhance high-frequency biases. Through these experiments shown in Fig. 16, we found that $\alpha_{1}=0.2$ and $\alpha_{2}=0.2$ works effectively.